### v3

# An Inquiry into Datacenter TCO for LLM Inference with FP8

Abstract.
As large language models (LLMs) continue to scale, their inference demands present significant challenges, particularly due to the high power consumption of AI accelerators in datacenters. These facilities require specialized cooling and power management systems, substantially increasing the total cost of ownership (TCO) for cloud service providers (CSPs). In this work, we analyze the computational characteristics and constraints of LLM inference from a TCO perspective, focusing on two representative accelerators: the Gaudi 2 and NVIDIA H100. We present a generalizable framework that enables CSPs to compare and select AI accelerators according to diverse operational requirements. Using this model, we analyze the impact of FP8 precision and LLM inference workload characteristics as key factors influencing TCO. We investigate FP8 quantization, which is gaining adoption in LLM training, as a technique to improve inference throughput while maintaining cost efficiency. Furthermore, our analysis of LLM inference workloads reveals that performance on thin GEMMs, which dominate the decode phase, can have a greater impact than theoretical hardware peak performance. By studying the interaction between power consumption, quantization strategies, and hardware architecture, we offer insights that support informed deployment decisions and guide future accelerator designs to improve the TCO of LLM inference.
LLM inference, TCO analysis, FP8 quantization
conference: Arxiv; 2025; Arxiv
1. Introduction
With the rapid advance of large language models (LLMs), AI datacenters have emerged as a critical infrastructure for supporting scalable and efficient inference workloads. The recent emergence of ”reasoning” AI models for both proprietary models such as OpenAI’s o1 (OpenAI, 2024) and Claude 3.7 (Anthropic, 2025), as well as open-source models such as DeepSeek R1 (DeepSeek-AI, 2025), has solidified the expectation that test-time scaling of LLMs will become a central workload for cloud service providers (CSPs). As LLM inference requests surge in both quantity and complexity, the volume of output tokens per request is also expected to grow substantially. This trend places significant pressure on datacenter infrastructure, making the selection of appropriate hardware a critical factor for sustaining performance and service-level efficiency. In this context, competition among AI accelerator manufacturers has intensified, with each aiming to position themselves to take advantage of the market.
Hardware selection has a direct impact on the total cost of ownership (TCO), which in turn influences the overall profitability and operational viability of AI datacenters. This necessitates a detailed examination of the various components that impact TCO, ultimately aiming to identify effective strategies for minimizing it. To provide practical guidance for CSPs, we construct an analytical model to compare hardware options, taking into account both facility costs and target performance levels. Although extensive research has explored the design of advanced AI accelerators (Micikevicius et al., 2022; Lee et al., 2025) and LLM quantization techniques (Lin et al., 2024; Ashkboos et al., 2024; Xiao et al., 2024), there remains a lack of rigorous investigation into how these developments impact the TCO for datacenter-scale LLM inference. Building upon our model, we analyze several key factors that influence TCO and discuss potential approaches for its optimization.
First, we analyze the impact of data precision in LLMs, with a particular focus on FP8 quantization, which has seen increasing adoption in recent state-of-the-art models such as DeepSeek v3 (DeepSeek-AI, 2024) and Llama 4 (Meta, 2025). Quantization offers the additional benefit that modern AI accelerators often include dedicated hardware units designed to accelerate low-precision computation, thereby improving throughput. In particular, FP8 has emerged as a promising datatype due to its lower computational overhead, reduced power consumption, and improved memory efficiency compared to commonly used formats such as FP32 and BrainFloat16 (BF16) (Kalamkar et al., 2019), especially for large batch sizes. With the adoption of FP8 by next-generation accelerators such as Intel’s Gaudi HPU and NVIDIA’s Hopper GPU, this format is gaining traction in cutting-edge LLMs. In this context, we explore how different FP8 quantization strategies impact both accuracy and throughput, examining their effects at two levels: matrix multiplication operations and end-to-end LLM inference.
Second, we investigate the behavior of LLM inference workloads when executed with FP8 quantization, aiming to understand its impact on throughput and resource utilization. To leverage quantization in a cost-effective manner, it is essential to understand the computational characteristics of LLM inference, which differ significantly from those of training. Even within the inference process, throughput varies across two distinct phases: the prefill phase and the decode phase. In particular, the decoding phase is dominated by Generalized Matrix-Vector (GEMV) operations and thin General Matrix-Matrix Multiplication (GEMM), both of which exhibit low computational intensity (CI) and lead to reduced Model FLOP Utilization (MFU) (Chowdhery et al., 2023; Narayanan et al., 2021). Furthermore, performance varies significantly across hardware platforms, depending on factors such as memory interface bandwidth and matrix computation unit throughput. Therefore, we conduct an in-depth analysis of the characteristics of the H100 and Gaudi2 in the context of LLM inference workloads. We further identify the performance limitations of each accelerator by examining their inherent architectural bottlenecks. To summarize our contributions:
  1. (1)
    We present a scalable TCO estimation model that serves as a generalizable framework for selecting optimal hardware configurations tailored to the needs of CSPs. While previous studies (Patel et al., 2024; Griggs et al., 2024) have analyzed throughput-cost trade-offs, they are often limited to specific deployment scenarios. In contrast, our model accommodates a wide range of cost structures and performance requirements, enabling flexible evaluation across different CSP contexts. Moreover, the model offers intuitive guidance for TCO optimization from a throughput-centric perspective.
  2. (2)
    Using this model, we investigate two key factors influencing TCO: FP8 numerical precision and LLM inference workload characteristics. Despite FP8’s growing popularity as a low-precision format that offers favorable trade-offs between throughput and memory efficiency, prior work has largely overlooked its empirical evaluation, particularly in the context of hardware-specific behavior during LLM inference. To address this gap, we conduct a comprehensive analysis of FP8 performance across different accelerator platforms. In addition, we provide a focused analysis of LLM decoding throughput that goes beyond a simple inspection of hardware specifications. Under FP8-quantized conditions, we further explore the performance characteristics and bottlenecks of each device within LLM workloads, offering insights that inform both hardware design and deployment strategies.
2. Exploring the TCO-throughput trade-off
People who buy and sell chips think about the cost of chips. People who operate data centers think about the cost of operations. … Our TCO is so good that even when the competitor’s chips are free, it’s not cheap enough.
Jensen Huang, Founder & CEO of NVIDIA
2.1. Understanding the TCO model
Figure 1. TCO comparison table for fixed traffic. For any desired task, given the throughput ratio between accelerators A and B $R_{Th}$ and server price ratio $R_{SC}$, the table shows the TCO of accelerator A as a proportion of the TCO of accelerator B, with $C_{S}=C_{I}$ assumed for clarity. A is therefore more cost-efficient than B when the obtained value is less than 1. The lower the value, the greater the cost-effectiveness of A.
When calculating the costs involved in running AI accelerators for datacenter operators, one must note that the purchase price of the chips themselves, while significant, constitute only part of the Total Cost of Ownership (TCO). For example, the purchase cost of the networking, cooling, and power infrastructure necessary to maintain the processors is frequently as large as the purchase cost of the chips and may even exceed it (Cui et al., 2017). As AI inference workloads constitute an ever greater portion of datacenter operations, achieving low cost and high efficiency requires careful planning. The choice of computing systems impacts both the throughput and the power consumption, both of which have a substantial influence on TCO.
Consider a scenario in which a CSP must choose between accelerators for a datacenter. In this scenario, two alternatives exist: system A, which is lighter-weight, delivers lower peak specifications, and has a lower purchase price, and system B, which provides substantially higher peak specifications but comes with a higher per accelerator purchase price. To identify which hardware option yields higher profitability, we must construct a model to compare TCO between the two hardware options.
When purchasing a server, it contains not only the chip but also networking, storage, and other associated components required for the server to function. Also, mulltiple chips may be located in a single server, whose costs must also be accounted for. In the following discussion, we denote $ServerCost$ as the term that combines the purchase costs of the multiple accelerator chips, CPUs, storage, and networking devices required to constitute a operational server.
$InfraCost$ indicates the infrastructure costs associated with operating the datacenter, including cooling and power provision, divided by the number of accelerators. We simplify the analysis by including operating expenses (OpEx), such as electricity consumption, in $InfraCost$. To understand the infrastructure cost, note that while the cost of electricity increases almost linearly with regard to the number of processors, the cost of the associated infrastructure, such as power racks and cooling equipment, is nearly constant. Because datacenter racks can only provide a limited amount of power, the per-chip cost of infrastructure is inversely proportional to the number of servers that can fit in a single rack. Therefore, the benefits of lower power consumption are twofold: the electricity consumption goes down and the number of processors per rack increases, spreading fixed infrastructure costs among a greater number of servers. For most datacenters, the cost of electricity per se is outweighed the cost of the rack and other equipment, even when amortized over several years, leading $InfraCost$ to be most heavily affected by the number of chips per rack.
We denote $R_{SC}$, $R_{IC}$, and $R_{Th}$ as the relative server costs, per server infrastructure costs, and throughputs between a single server of computing systems A and B, respectively, and define them as follows:
$R_{SC}=\frac{ServerCost_{A}}{ServerCost_{B}},\quad R_{IC}=\frac{InfraCost_{A}}{InfraCost_{B}},\quad R_{Th}=\frac{Throughput_{A}}{Throughput_{B}}$
Regardless of the devices in use, the amount of traffic remains constant. We therefore make an iso-traffic assumption and use $N$ to represent the number of B servers required to process the expected traffic. We use $C_{S}=ServerCost_{B}$ and $C_{I}=InfraCost_{B}$ to represent the cost components attributed to the server and the infrastructure cost per server when using system B. Making the assumption that the cost of infrastructure is the same for systems A and B, i.e., $R_{IC}=1$, derive the following equation:
(1) $\frac{TCO_{A}}{TCO_{B}}=\frac{C_{S}\times R_{SC}\times\frac{N}{R_{Th}}+C_{I}\times\frac{N}{R_{Th}}}{C_{S}\times N+C_{I}\times N}$
The numerator captures the TCO incurred when using the lower-performance system A to meet the same throughput target as system B, while the denominator reflects the baseline TCO of using system B for a given amount of traffic. As total traffic remains constant regardless of the underlying device, purchasing system A will require $R_{Th}/N$ nodes, leading to possibly increased infrastructure costs, depending on the throughput ratio $R_{Th}$. At the same time, its lower per-chip cost, reflected in the server price ratio $R_{SC}$, could result in a lower TCO.
Based on this model, Figure 1 illustrates the TCO comparison between two computing systems, expressed as an $A/B$ ratio. The x-axis represents the server price ratio, and the y-axis represents the per server throughput ratio of the two systems for the desired task. Since exact costs are often confidential in real-world data centers, expressing TCO in relative ratios allows for flexible adaptation to various operational contexts. In deriving the values in Figure 1, we assumed $C_{S}=C_{I}$, assuming that server and infrastructure costs contribute equally (i.e., 50% each) to the overall TCO, thereby allowing these terms to cancel out across the numerator and denominator. We also assume that both systems consume the same power per server and that the same number of servers fit in a single rack, i.e., $R_{IC}=1$. In practice, however, it may be appropriate to substitute variables to reflect each specific deployment scenario. The values in Figure 1 are an example of the TCO ratio between two hardware options. The lower the value, the less expensive, and hence more cost-effective, system A becomes and the more appealing it is to adopt. Conversely, a ratio above 1.0 suggests that system B is preferable from a TCO perspective.
2.2. TCO factor analysis
The TCO analysis presented so far offers a flexible and adaptable methodology for evaluating system options across diverse data center environments. In the implemented TCO model, both the server price ratio and throughput ratio can vary significantly depending on factors such as server acquisition costs, accuracy constraints, and the specific tasks being executed. The cost of processors may change depending on negotiations with manufacturers and are generally confidential.
A key contribution of the analysis is that it can incorporate throughput ratios of any desired task, e.g., LLM decoding as opposed to LLM training or prefill, allowing users to take the intricacies of hardware-software co-optimization into account. For instance, the service may require that only models that pass a quality assurance (QA) review are allowed, and that the model meets specific latency requirements when deployed. All of these factors can be incorporated into the TCO model to provide insights. For example, the prefill and decode phases of LLM inference exhibit markedly different throughput profiles. As shown in Patel et al. (2024), it is possible to use different chips and configurations for the prefill and decode to optimize TCO as their relative throughputs for the prefill and decode tasks are different. When these different throughputs are compared using our TCO model, it becomes easier to grasp why it is profitable to utilize different hardware and configurations for different inference phases.
To build an efficient datacenter environment, CSPs must consider complex trade-offs between functionality, performance, and cost. Quantization has emerged as a viable solution to overcome some of these limitations by reducing memory usage. Leveraging the inherent error resilience of LLMs enables reduced bit-width representations of model parameters while maintaining comparable output quality. Consequently, incorporating these lower-precision modes into TCO analysis becomes essential. In the following section, we explore FP8 quantization as a particularly important technique.
3. FP8 for high-throughput LLM inference
In this section, we provide an overview of FP8 as a numerical format, analyze how it is implemented across hardware backends, and evaluate its performance in GEMM operations, one of the core building blocks of LLM inference.
3.1. Motivation for FP8 precision
While BF16 precision is commonly used in recent LLMs, there is substantial opportunity to boost throughput and to minimize TCO by operating at a lower precision level, FP8, within acceptable bounds of accuracy degradation. Especially with the emergence of next-generation foundation models such as DeepSeek v3 and LLaMA 4 whose parameters were trained in FP8 precision, the FP8 GEMM has become increasingly central to achieving high throughput at scale. Moreover, even for models originally trained using BF16, post-training quantization (PTQ) techniques now enable effective conversion to FP8 without significant accuracy degradation. However, the effectiveness of low-precision computation is highly dependent on the underlying hardware architecture. As precision decreases, achieving full hardware utilization becomes more challenging, directly affecting throughput. Therefore, hardware-specific characteristics are critical in determining actual performance at lower precision levels. In this section, we explore the characteristics and implications of Llama 4 style FP8 precision, analyzing its computational behavior and potential for performance optimization.
3.2. Differences in hardware capabilities
We compare implementations of FP8 GEMM between NVIDIA GPUs and Intel Gaudi HPUs, identifying key differences despite both adhering to the FP8 specification (Micikevicius et al., 2022).
(Accumulation precision) Hopper GPUs use a 14-bit accumulator for FP8 GEMMs (DeepSeek-AI, 2024), requiring casting to CUDA cores for higher precision. Software optimizations, such as applying FP32 accumulation to only one in four WGMMA instructions, reduce error but increase kernel complexity and remain less precise than full FP32 accumulation. In contrast, Gaudi HPUs always accumulate in FP32 (Lee et al., 2025), ensuring higher numerical precision.
(E4M3 range) The Gaudi 2 follows the IEEE specification for special values, unlike NVIDIA GPUs, which use a single special value representation (Noune et al., 2022). This results in seven fewer magnitude representations and a maximum value of 240 for E4M3 in the Gaudi 2 compared to 448 on NVIDIA GPUs. This has been addressed in the Gaudi 3, but we were unable to test these in our experiments.
(Power-of-2 scaling) Gaudi HPUs allow the modification of floating-point exponent biases to accelerate scaling factor application. The Gaudi 2 supports fixed hardware-accelerated scaling factors of $2^{-8},2^{-4},2^{0},2^{4}$ for E4M3, while the Gaudi 3 extends this to arbitrary powers of 2 between $2^{-32}$ and $2^{31}$. However, this feature is limited to GEMMs with per-tensor scaling factors.
(Stochastic rounding) During FP8 quantization, stochastic rounding can be applied when converting 16/32-bit floating-point values to FP8, similar to the technique proposed for FP32-to-BF16 conversion in (Fan et al., 2024). This method is distinct from stochastic rounding applied at the inner-product (El Arar et al., 2023).
(Sparsity) NVIDIA GPUs support semi-structured sparsity acceleration, potentially doubling tensor core peak throughput. However, despite attempts to leverage sparsity in LLMs (Frantar and Alistarh, 2023; Sun et al., 2024), dense GEMMs remain dominant due to accuracy loss and limited speedups. Gaudi HPUs do not support sparsity acceleration.
Table 1. Throughput and power measurements for square FP8 GEMMs with row-wise scaling. Ratios of measured TFLOPS and power to their peak values are shown in parentheses. H100 has peak FP8 throughput of 1989.9 TFLOPS with a TDP of 700W, and Gaudi 2 has 865 TFLOPS with a TDP of 600W.
Device (M,K,N) TFLOPS Power (W) TFLOPS/W
Gaudi 2 1K 367.9 (42.5%) 375 (63%) 1.0
2K 586.2 (67.8%) 460 (77%) 1.3
4K 817.1 (94.5%) 460 (77%) 1.8
8K 741.8 (85.8%) 490 (82%) 1.5
H100 1K 218.3 (11.0%) 350 (50%) 0.6
2K 879.7 (44.2%) 690 (99%) 1.3
4K 1167.6 (58.7%) 690 (99%) 1.7
8K 1084.7 (54.5%) 690 (99%) 1.6
3.3. GEMM throughput measurements
For a GEMM between matrices of size $(M\times K)\times(K\times N)$, the total number of floating-point operations (FLOPs) performed is $2MKN$. This is derived from the $M\times N$ dot products of length $K$, where each element undergoes one multiplication and one addition. Following the convention where FLOPS denotes FLOPs per second, we compute throughput in FLOPS using the theoretical FLOPs and the measured latency. We can thus calculate the MFU by dividing observed throughput by peak throughput.
As LLM computations are dominated by matrix multiplications, GEMM throughput serves as an upper bound for model end-to-end performance. While accelerator specifications list peak throughput values, actual throughput rarely reaches this limit, particularly for small matrices. Table 1 presents GEMM throughput and power consumption measurements for row-wise scaled FP8 GEMM. We conducted measurements on NVIDIA H100 GPUs using the NGC PyTorch 25.03 image and on Intel Gaudi 2 HPUs using the Synapse AI v1.12.0 image. Power consumption was recorded via the “nvidia-smi” and “hl-smi” utilities, respectively. Input casting overheads were excluded from the measurements. Additional results for GEMM throughputs under different conditions are provided in Tables 2 and 3. Our results show that the Gaudi 2 achieves higher utilization than the H100, particularly for small matrices. The performance drop for smaller matrices is steeper on the H100, with the Gaudi 2 providing a higher TFLOPS at matrix sizes of 1K. Also, the Gaudi 2 exhibits lower power consumption than its stated 600W TDP, whereas the H100 approaches its peak 700W TDP even at moderate utilizations.
Unlike NVIDIA GPUs, which incorporate multiple small matrix multiplication accelerators, Gaudi HPUs feature a few large matrix accelerators. As a result, Gaudi HPUs requires fewer input elements per cycle to fully utilize compute resources, thereby reducing first-level memory bandwidth requirements and improving efficiency (Intel, 2022, 2024). Additionally, Gaudi HPUs leverage a graph compiler to dynamically reconfigure the MME size based on the target GEMM dimensions, optimizing resource utilization. This superior efficiency of the Gaudi 2 for small matrices is particularly relevant given that LLM hidden dimensions typically range between 1K (e.g., Llama 1B) and 8K (e.g., Llama 70B), with tensor parallelism further reducing matrix sizes. Moreover, the lower power draw of the Gaudi 2 relative to its TDP, as shown in Table 1, suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
Table 2. Gaudi 2 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead. Measured throughput relative to the peak FP8 throughput (865 TFLOPS) is included in parentheses. Hardware acceleration is only available for per-tensor scaling.
Type Size Per-Row Per-Tensor HW Accel.
E4M3 1K 494 (57.1%) 494 (57.1%) 494 (57.1%)
2K 506 (58.5%) 641 (74.1%) 641 (74.2%)
4K 735 (84.9%) 796 (92.1%) 801 (92.6%)
8K 742 (85.7%) 822 (95.0%) 852 (98.4%)
E5M2 1K 306 (35.4%) 494 (57.1%) 493 (57.0%)
2K 506 (58.5%) 642 (74.2%) 642 (74.2%)
4K 735 (84.9%) 802 (92.7%) 802 (92.7%)
8K 726 (83.9%) 825 (95.4%) 825 (95.4%)
Table 3. H100 throughput (TFLOPS) for scaled FP8 GEMMs on square matrices of the given size, excluding quantization overhead. Ratios relative to peak FP8 throughput are shown in parentheses. Only E4M3 results are reported, as Python functions for E5M2 GEMM are unavailable for GPUs.
Accum. Size Per-Row Per-Tensor
FP32 1K 217 (11.0%) 186   (9.4%)
2K 299 (15.1%) 840 (42.4%)
4K 362 (18.3%) 1099 (55.5%)
8K 396 (20.0%) 1300 (65.7%)
Fast 1K 237 (12.0%) 147   (7.4%)
2K 810 (40.9%) 896 (45.3%)
4K 1136 (57.4%) 1205 (60.9%)
8K 1123 (56.8%) 1388 (70.1%)
4. FP8 configuration impact on accuracy
Table 4. The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration. The “Measured” column includes results for dynamic “FP8 (D)” scaling. BF16 accuracies are included for both columns to control for evaluation condition differences.
Llama v3.1 8B Inst. Cited Measured
BF16 FP8 (S) BF16 FP8 (D)
MMLU CoT 5-shot 68.4% 66.3% 68.8% 68.3%
GSM8K CoT 5-shot 75.7% 70.4% 83.1% 84.5%
Winogrande 78.1% 77.6% 73.8% 73.8%
TruthfulQA mc1 37.1% 34.9% 39.9% 39.4%
TruthfulQA mc2 54.0% 52.2% 55.1% 54.3%
In this section, we investigate how different FP8 quantization configurations affect the accuracy of large language models. Specifically, we evaluate the impact of scaling strategies (static vs. dynamic), format types (E4M3 vs. E5M2), and rounding methods (stochastic vs. deterministic). Our experiments are conducted on instruction-tuned variants of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B. Unless otherwise specified, dynamic row-wise scaling is applied to all linear layers, while the LM head remains in BF16.
4.1. Dynamic vs static scaling
We examine the trade-offs between static and dynamic scaling strategies for activation quantization. Static scaling offers higher throughput for large matrices as scaling factors do not need to be computed for each input. It also enables per-tensor quantization, leading to improved GEMM utilization. In contrast, dynamic scaling enhances output quality by assigning separate scaling factors to each token. Native FP8 foundation models such as DeepSeek v3 and Llama 4 use dynamic scaling for both training and inference.
Table 4 presents a comparative analysis of model output quality under static and dynamic FP8 scaling. The results show that dynamic quantization achieves accuracy comparable to the original BF16 models, while static quantization often results in noticeable degradation. Although various techniques, such as smoothing (Xiao et al., 2024) and rotation-based methods (Liu et al., 2024; Ashkboos et al., 2024), can help mitigate this degradation, they rely on calibration data and may yield suboptimal scaling factors for unseen inputs. As a result, dynamic quantization remains the preferred approach, assuming the associated throughput overhead is within acceptable bounds.
Table 5. Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models. Stochastic rounding provides little or no benefit to accuracy while E5M2 is detrimental, especially for smaller models. Experiments were conducted on a Gaudi 2 HPU.
Model Data Type Rounding MMLU
Llama v3.2 BF16 - 46.3%
1B Instruct E4M3 SR 45.7%
E4M3 RTN 45.5%
E5M2 RTN 44.5%
Llama v3.2 BF16 - 61.8%
3B Instruct E4M3 SR 61.7%
E4M3 RTN 61.6%
E5M2 RTN 60.7%
Llama v3.1 BF16 - 68.8%
8B Instruct E4M3 SR 68.3%
E4M3 RTN 68.3%
E5M2 RTN 67.5%
Llama v3.3 BF16 - 82.0%
70B Instruct E4M3 SR 82.0%
E4M3 RTN 82.0%
E5M2 RTN 82.2%
4.2. E4M3 vs. E5M2
Prior work has shown that E4M3 yields better accuracy than E5M2 on language tasks (Shen et al., 2024). We extend this analysis to instruction-tuned models, as summarized in Table 5. To identify the optimal format for inference, we compare the two using MMLU accuracy measured with LM Evaluation Harness (v0.4.7 (Gao et al., 2024)). Across all evaluated models, E4M3 consistently outperforms E5M2. Moreover, Table 2 shows that GEMM throughputs for both formats are comparable, making E4M3 the preferred choice despite its smaller representational range on the Gaudi 2.
Figure 2. Process and utilization characterization of the two phases of generative LLM inference.
4.3. Stochastic rounding
Hardware-accelerated stochastic rounding for FP8 quantization is a unique feature of Gaudi HPUs, inspired by techniques such as those proposed in (Fan et al., 2024), and is not supported on NVIDIA GPUs. Equation 2 formalizes the concept, where a higher-precision value $x$ is rounded up to $x_{up}$ or down to $x_{down}$ stochastically based on the distance from $x$.
(2) $x_{quantized}=\begin{cases}x_{up}\quad(p=\frac{x-x_{down}}{x_{up}-x_{down}})\\ x_{down}\quad(p=\frac{x_{up}-x}{x_{up}-x_{down}}).\end{cases}$
Stochastic rounding is expected to preserve more of the original numerical information post-quantization, potentially leading to improved model accuracy. However, empirical results in Table 5 indicate that stochastic rounding during quantization does not significantly enhance model accuracy. Furthermore, it may even lead to accuracy degradation in certain cases (Fan et al., 2024). These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remain inconclusive.
5. Phase-aware analysis of FP8 LLM inference
LLM inference comprises two distinct computational phases, prefill and decode, each exhibiting different performance characteristics and hardware bottlenecks. In this section, we analyze how FP8 quantization affects each phase, focusing on the interaction between matrix shapes, computational intensity, and hardware utilization.
Figure 3. Comparison of measured throughput and the roofline throughputs obtained from hardware specifications. While the H100 offers significantly higher theoretical throughput, Gaudi 2 achieves superior measured performance in many decoding settings. Measurements were conducted using a 400W power cap per chip for fair comparison.
5.1. LLM inference phases
Generative LLM inference comprises two distinct phases: a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024). Figure 2 illustrates the key differences between these phases. During the prefill phase, the LLM generates the first token based on the provided input prompt. Throughout this process, key-value (KV) pairs are stored in the KV cache for reuse in subsequent token generations. The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound. This high degree of parallelism allows for effective hardware utilization, enabling efficient input processing.
Once the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens. Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV and thin GEMM operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase. Modern inference frameworks such as vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023) partially remedy this issue by batching multiple decoding requests, improving the computational intensity of the linear layers. However, the batch size is limited by the memory capacity as each sequence in a batch requires its own KV cache. Furthermore, the computational intensity of the attention operation during decoding is unaffected by batch size, instead requiring a GEMV for each sequence. This slowdown becomes more evident at longer sequence lengths, where the attention FLOPs continue to increase in proportion to sequence length while the FLOPs required for linear layers remain constant. Grouped query attention (GQA) (Ainslie et al., 2023) converts the operation into a thin GEMM, but it remains memory-bound. Multi-head latent attention (MLA) (DeepSeek-AI et al., 2024) further improves the computational intensity during the decode phase, but the trend remains the same.
5.2. Calculating inference FLOPs
While previous works (Patel et al., 2024) have evaluated inference performance using time to first token (TTFT) and time per output token (TPOT), these metrics do not facilitate comparisons across different inference stages, model sizes, or sequence lengths. To ensure a consistent comparison, we directly compute model FLOPs. We follow the methodology described in (Ansel et al., 2024; Narayanan et al., 2021), computing only the FLOPs associated with matrix multiplications and excluding those related to autoregressive attention masking, which can often be skipped in practice (Dao, 2024). This yields a more faithful representation of the actual compute cost when the KV cache is utilized for incremental decoding.
Using this method, the FLOPs required for a forward pass of a Llama model with $l$ transformer blocks, hidden size $h$, intermediate size $ah$, head size $d$, head count $H=h/d$, GQA group size $g$, vocabulary size $v$, and input sequence of length $s$ can be calculated as follows:
(3) $f_{llama}(s)=2sh^{2}l(3a+2+\frac{2}{g})+2s^{2}hl+2vsh.$
By denoting $A=(3a+2+\frac{2}{g})$ as a model-specific constant, Equation 3 simplifies to:
(4) $f_{llama}(s)=2s(Ah^{2}l+vh)+2s^{2}hl.$
When the model generates $t$ tokens in a single decoding iteration where $t\ll s$, we approximate the additional computation by evaluating Equation 3 at $s^{\prime}=s+t$, yielding:
(5) $f_{llama}(s+t)-f_{llama}(s)$
$\approx 2t(Ah^{2}l+vh)+4sthl.$
From Equation 5, we observe that the computational cost of the linear layers, including the LM head, remains independent of the previous sequence length $s$, while the attention-related FLOPs scale proportionally with $s$. In the autoregressive setting where $t=1$, and for a batch of $b$ sequences with respective sequence lengths $s_{1},\dots,s_{b}$, the FLOPs required for one decoding step can be approximated as:
(6) $2b(Ah^{2}l+vh)+4hl\sum_{i=1}^{b}s_{i}.$
A key observation is that only the first term, $2bAh^{2}l$, corresponding to the linear layers (excluding the LM head), is computed in FP8. In contrast, the $2bvh$ term for the LM head and the $4hl\sum_{i=1}^{b}s_{i}$ term for attention are computed in BF16. Additionally, online dequantization of the KV cache introduces extra overhead not accounted for in Equation 6. However, these overheads do not contribute to the model FLOPs as defined in the MFU metric (Chowdhery et al., 2023).
A more fundamental limitation is that not every FLOP can be executed at full efficiency due to hardware utilization constraints. For instance, the Gaudi 2 features a peak HBM bandwidth of 2.4 TB/s and a peak FP8 GEMM throughput of 865 TFLOPS, which implies that a FLOP/byte ratio (i.e., computational intensity, CI) of at least 360 is required to saturate the compute units. However, during the decoding phase, the workload often reduces to a thin GEMM of shape $(b\times h)\times(h\times ah)$, where $b\ll h$ and $a\geq 1$. In this case, the resulting CI drops to approximately $2b$ for FP8 and $b$ for BF16, which is far below the required threshold for achieving peak throughput at practical batch sizes. Moreover, matrix multiplication units typically operate on fixed-size blocks, and full hardware utilization is only achieved when input dimensions align with hardware-friendly shapes, such as multiples of 128 (Lee et al., 2024).
Furthermore, KV cache computations introduce a distinct bottleneck. Increasing the batch size does not improve CI, as each sequence in the batch maintains an independent KV cache. In the case of a BF16 KV cache with GQA using $g$ groups, the CI is bounded by $g$ FLOPs/byte. As a result, even with an optimally implemented kernel, the theoretical throughput for applying the query to the KV cache is limited by the product of memory bandwidth and CI. For example, in LLaMA v3 models with $g=8$ on a Gaudi 2 with 2.4 TB/s peak HBM bandwidth, the maximum achievable throughput is 19.2 TFLOPS, only a small fraction of the device’s peak GEMM throughput. Because attention operations are inherently memory-bound and their FLOPs scale linearly with sequence length, the decoding phase at longer sequence lengths eventually becomes limited by attention throughput. Understanding the distinct computational characteristics of each phase allows for more targeted optimization strategies, particularly in improving the efficiency of the decode phase, which dominates end-to-end inference latency.
5.3. Prefill phase
We first conducted an analysis of the operational characteristics of the two accelerators in the prefill phase. Figure 4 presents a roofline comparison of prefill throughput (TFLOPS) for three models across varying sequence lengths on the H100 and Gaudi 2. The results show that the H100 consistently achieves significantly higher throughput than the Gaudi 2 in this phase, reaching up to double the performance of the Gaudi 2 for LLaMA 8B.
In the prefill phase, attention layers process the entire input sequence at once, enabling large matrix multiplications to be executed in a single step. This leads to high utilization of compute engines and results in performance primarily governed by GEMM throughput, rather than memory bandwidth or utilization constraints.
As model size increases, matrix dimensions scale proportionally with the hidden size, making computations increasingly compute-bound. This results in a clear trend of improved prefill throughput for larger models, highlighting the importance of raw compute performance in this phase. A similar trend is observed with longer input sequences, where hardware utilization improves until saturation is reached.
Figure 4. Roofline diagram during the prefill phase for batch size 1. Throughput improves with longer sequence lengths until they begin to decline as the proportion of attention computations, which are slower than GEMMs, takes up a greater share of the computation. We use static FP8 scaling for both the H100 and the Gaudi 2 as optimized dynamic FP8 scaling is unavailable for the Gaudi.
Figure 5. Decode throughput comparison between BF16 and FP8 on Llama v3.1 8B Instruct using a batch size of 64 using static FP8 scaling on the Gaudi 2 (left); static and dynamic FP8 scaling on the H100 (right). Throughput differences between FP8 and BF16 in the Gaudi 2 are 50% or greater, whereas they are under 25% for the H100. Dynamic scaling has higher throughput despite its greater overhead because row-wise GEMM is faster than per-tensor GEMM for small matrices on the H100 (see Table 3).
5.4. Decode phase
The decode phase has emerged as the primary contributor to overall inference latency in modern LLMs. As “reasoning” models (DeepSeek-AI, 2025) often generate thousands of tokens to improve output quality, end-to-end performance is increasingly dominated by the decode phase. Consequently, optimizing this phase is critical to reduce inference latency and maximize throughput.
Figure 5 compares decode throughputs on the H100 across three settings: BF16, dynamic FP8 scaling, and static FP8 scaling. On the H100, FP8 yields less than a 25% improvement over BF16, whereas the Gaudi 2 achieves up to a 50% gain. These results indicate that the H100 does not fully realize the theoretical FP8 speedup (e.g., double that of BF16) during the decode phase. As a result, its FP8 performance is only marginally better than BF16. Notably, Gaudi 2 with FP8 achieves comparable decode throughput to the H100, despite having significantly lower peak GEMM throughput.
5.5. Power capping strategies
As generative LLM inference tasks become more prevalent, datacenters must contend with substantial power consumption from AI accelerators, leading to high TCO. High-performance GPUs can draw significant power, but data centers have limited power infrastructure, making efficient power management crucial. One promising strategy is GPU power capping, which imposes a fixed power limit for GPUs. This keeps overall power usage predictable and helps prevent out-of-spec conditions, such as rack overheating or hardware damage.
To observe the performance of each GPU in constrained conditions, we measured the throughput of the Gaudi 2 and H100 in a constrained power environment. Given that much existing infrastructure in production datacenters has been built around the A100, with infrastructure commonly optimized for its 400W TDP, we perform our evaluations under the same power budget to ensure a fair and compatible comparison. As noted in Patel et al. (2024), the decode phase does not show throughput deterioration with power capping at 400W. Figure 3 compares the theoretical roofline with the measured TFLOPS under a 400W power constraint. We show only the 400W measurements as the measurements without power capping are nearly indistinguishable during decoding. Note that while the H100 exhibits a much higher theoretical throughput, the Gaudi 2 achieves higher measured performance in several scenarios.
However, both devices currently support per-GPU power capping only, rather than per-rack power capping. Per-rack power capping offers a compelling solution by preserving the benefits of power management while allowing each GPU greater flexibility in power consumption. When power is constrained at the individual GPU level, no device can exceed its fixed limit, potentially leaving unused power headroom untapped. In contrast, per-rack capping treats the total power budget of each rack as a collective resource, letting one GPU utilize surplus capacity if another GPU draws less than its share. This approach can lead to more efficient resource usage and higher overall performance.
5.6. Thin GEMM with FP8 quantization
Figure 6. Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8. The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100. The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 6.
Thin GEMM is a critical workload that significantly impacts overall inference latency. Although LLM sizes continue to scale, their hidden dimensions grow at a much slower rate. For example, Llama 8B has a hidden size of 4096, whereas Llama v3.3 70B increases it only to 8192. As shown in Equation 6, the decoding stage at short sequence lengths ($s<h$) is dominated by linear operations. The growing adoption of fine-grained Mixture-of-Experts (MoE) models (DeepSeek-AI, 2024; Meta, 2025) further contributes to this trend, as a larger number of experts reduces the average number of activations assigned to each expert during batched decoding. Additionally, tensor parallelism (TP), which is commonly employed to reduce latency and memory overhead, further partitions model weight matrices into smaller submatrices. As a result, the efficiency of small matrix operations has become a key factor in optimizing LLM inference performance.
Unlike large matrix-matrix multiplications, thin GEMMs fail to fully utilize compute units, making it difficult to achieve peak performance. Moreover, the inefficiency that derives from hardware utilization increases significantly as low data precision is adopted. Theoretically, FP8 offers up to double the throughput of BF16 or FP16. However, even if the compute units themselves are capable of fast execution, suboptimal input/output pipelining or memory access patterns can severely bottleneck overall throughput. Therefore, to operate datacenters efficiently, it is crucial to understand how specific hardware platforms perform in scenarios involving thin GEMM and low-precision computation.
Based on the insight about the thin GEMM utilization difficulties, we examine the behavior of the Gaudi 2 and H100 under a combination of thin GEMM workloads and FP8 precisions. Table 6 presents the GEMM TFLOPS in a setting where the dimension of the matrix $M$ is significantly smaller than $K$ and $N$, which expresses the thin GEMM condition. Compared to its performance on large GEMMs, the H100 shows notably lower throughput than its specifications under this thin GEMM scenario. FP8 and BF16 throughputs for the H100 are nearly identical, failing to reflect the expected performance gains from lower precision. In contrast, the Gaudi 2 exhibits an FP8 throughput approximately double that of BF16, which aligns closely with theoretical expectations.
Table 6. Thin GEMM throughputs (in TFLOPS) for the Gaudi 2 and H100 measuring $(M\times K)\times(K\times N)$ GEMMs, where $M$ corresponds to typical batch sizes during inference and $K,N$ represent hidden dimension sizes. Row-wise scaling is used for FP8 GEMMs. Throughput scales linearly with $M$ on both devices, but the Gaudi 2 consistently outperforms the H100, even for BF16. Results highlight the superior efficiency of the Gaudi 2 for thin GEMMs, crucial for LLM decoding.
GEMM TFLOPS Gaudi 2 H100
Shape: (M,K,N) BF16 FP8 BF16 FP8
( 8, 1024, 1024) 3.3 3.8 1.7 1.7
(16, 1024, 1024) 6.5 11.4 3.4 3.9
(32, 1024, 1024) 12.8 23.8 6.5 7.0
(64, 1024, 1024) 26.7 54.0 12.6 14.9
( 8, 2048, 2048) 12.4 26.1 6.7 7.5
(16, 2048, 2048) 20.6 48.6 12.9 15.0
(32, 2048, 2048) 48.0 87.6 27.1 28.2
(64, 2048, 2048) 91.3 163.2 52.3 60.5
( 8, 4096, 4096) 18.8 35.4 14.4 16.8
(16, 4096, 4096) 37.4 67.9 28.6 33.5
(32, 4096, 4096) 73.6 132.0 68.3 68.1
(64, 4096, 4096) 144.5 253.4 133.3 133.9
These differences in FP8 thin GEMM performance are likely attributed to the architectural differences of the hardware platforms. In the case of NVIDIA Hopper, details of the Tensor Core architecture remain undisclosed, making it difficult to analyze the actual implementation. We expect that future disclosure of internal implementation details would enable a more systematic investigation. In contrast, Gaudi2 provides public documentation that contains their Matrix Multiplication Engine (MME) architecture, which is an output-stationary systolic array.
Figure 7 depicts the differences between the MME architectures of the Gaudi2 and those of other commercial hardware solutions. Unlike other commercial hardware offerings that employ multiple small compute units, Gaudi2 integrates a single large MME. In architectures that rely on several small units, each unit consumes input data in every cycle, incurring significant overhead in unit-to-unit data movement. In contrast, one large MME consumes a much smaller amount of input data per cycle and requires less expensive ALU-to-ALU communication within a whole MME. The effect of communication cost leads to notable MFU differences, as shown in Figure 6. The data movement latency is effectively hidden in compute-bound conditions, but it becomes critical in memory-bound thin GEMM operations, where communication overhead can dominate performance and markedly degrade overall performance.
Furthermore, Figure 8 illustrates the reconfigurable MME sizes of Gaudi 2. While the Gaudi2 architecture contains two MMEs with size $256\times 256$, they can be folded and integrated in different sizes with a smallest width 128. The size of the MME is optimized at run time via the Gaudi Graph Compiler. This flexibility allows the hardware to adapt to a wide range of GEMM dimensions, offering substantial potential for improving LLM inference efficiency.
5.7. Exponential function optimization
While the Gaudi 2 outperforms the H100 in the LLM decode phase for short sequence lengths, its throughput gains diminish at longer sequence lengths. We find that this decline is primarily due to the absence of the dedicated Special Function Units (SFUs, also referred to as multi-function units, or MUFUs) in the Gaudi 2 & 3 to accelerate the exponential calculations in the softmax.
During typical LLM inference, SFUs perform crucial tasks. Although GEMM operations that leverage tensor cores dominate LLM workloads, specialized functions such as softmax and SiLU are also necessary. For decoding workloads, the ratio of exponential computations relative to GEMM computations increase in proportion to the sequence length. As a result, the softmax operation becomes a bottleneck for decoding at long sequence lengths. GPUs equipped with SFUs can process exponential functions on SFUs while simultaneously executing GEMMs on tensor cores, effectively parallelizing the workload and boosting overall performance.
In contrast, both the Gaudi 2 & 3 have a disadvantage in that they lack SFUs and must handle exponential computations via their TPC cores. They also have a lower ratio of vector core compute relative to matrix computation units compared to the H100. Peak TPC core BF16 throughput is 11 TFLOPS for the Gaudi 2 and 28.7 TFLOPS for the Gaudi 3, whereas the H100 has 133.8 TFLOPS BF16 CUDA core throughput and includes separate SFU cores. This implies that despite the notable efficiency of thin GEMMs in the Gaudi 2, exponential function calculations emerge as the main bottleneck for long sequences, as the softmax computation scales $O(BS)$ for batch size $B$ and sequence length $S$ during decoding.
The Gaudi 2 incorporates an embedded look-up table (LUT), which could theoretically store precomputed special-function values via piecewise linear approximation. If the LUT offers sufficiently low access latency and can hold a large enough range of values, it may be a valid substitute for SFUs. However, Gaudi 2 currently lacks software support for programming the LUT. Although it has strong potential to improve performance with long sequences, the limited software stack constrains its practical effectiveness.
Figure 7. Gaudi2 employs one large systolic array unit for each MME, unlike the many other commercial hardware employs many small units.
Figure 8. Dynamically reconfigurable MME architecture on Gaudi for target GEMM optimization.
6. Throughput shifts the TCO balance
As described in Section 2, the TCO varies significantly depending on factors such as data precision, workload characteristics, and batch size. Figure 9 illustrates how the TCO ratio changes under varying conditions, particularly focusing on two key parameters: the throughput ratio and the server cost ratio. In our TCO model, system A represents the Gaudi 2 and system B represents the H100. In Figure 9, movement along the x-axis reflects the server cost ratio, which is determined by factors such as device procurement cost, infrastructure constraints, and other datacenter-specific considerations. This ratio varies across CSPs depending on their individual deployment environments. Therefore, the dominant factor influencing the TCO ratio is the throughput ratio ($R_{Th}$).
In the LLM decoding phase, which dominates overall inference latency, the Gaudi 2 achieves nearly 150% the throughput of BF16 when using FP8 quantization, while the H100 shows little to no performance gain. As a result, $R_{Th}$ increases, and the TCO ratio shifts upward along the y-axis, approaching the green region in the figure, indicating higher cost efficiency for the Gaudi 2. Conversely, in scenarios involving long sequences, the Gaudi 2 suffers from the softmax bottleneck due to its lack of a dedicated SFU or an alternative workaround. This lowers the throughput ratio, shifting the TCO ratio downward along the y-axis toward the red region, which implies reduced cost efficiency for the Gaudi 2.
Ultimately, the actual position on the TCO map depends on the specific conditions of each datacenter and on the specific operations required by the CSP. Therefore, accurate evaluation requires a detailed analysis of the operational environment and workload profile unique to each CSP.
Figure 9. Visualization of TCO Variation by Server Cost and Throughput Ratios.
7. Conclusion
In this study, we present a scalable framework for comparing AI accelerator options in terms of datacenter TCO. Using our model, we examine two key factors that influence throughput and are directly connected to the TCO: numerical precision and LLM inference workload characteristics. We evaluate two representative next-generation accelerators, the NVIDIA H100 and Intel Gaudi 2, and analyze their behavior under FP8 quantization.
However, our study has several limitations. First, our analysis is based on a case study focused specifically on Llama-based models in the context of inference. As such, the results may not directly generalize to other workloads, such as training. Second, we adopt FP8 precision to reflect anticipated trends in future-generation models. However, the applicability of our findings depends on the extent of hardware support for FP8, which remains a limitation.
While our study provides valuable insights into cost-performance trade-offs, additional research is required to generalize the results to full-scale datacenter environments with diverse infrastructure, workload mixes, and deployment strategies. We hope that the insights provided in this study help inform deployment decisions and guide future accelerator design efforts aimed at improving the TCO of LLM inference in datacenter environments.